# D. Leykekhman, B. Vexler and J. Wagner

### v1

@@ 4.1 Spatial discretization
This assumption is valid, for example, for Taylor-Hood and MINI finite elements on shape regular meshes, see [3, Assumption 7.2]. We introduce the vectorial discrete Laplace operator $-\Delta_{h}\colon\mathbf{X}_{h}\rightarrow\mathbf{X}_{h}$ by
$(-\Delta_{h}{\bf{v}}_{h},\chi)=(\nabla{\bf{v}}_{h},\nabla\chi),\quad\forall{\bf{v}}_{h},\chi\in\mathbf{X}_{h}.$ (4.2)
We also define the space of discretely divergence-free vector fields $\mathbf{V}_{h}$ as
@@
Note that, if ${\bf{u}}\in\mathbf{V}$, then it holds $R_{h}^{S}({\bf{u}},p)\in\mathbf{V}_{h}$. Further, the Stokes Ritz projection satisfies the following stability, see [4, Theorem 5.2.1].
@@
Let us recall the following error estimates for the stationary discrete Stokes problem: See [4, Theorem 5.25], [22, Theorems 4.21, 4.25, 4.28], [15, Theorems 53.17 & 53.19] or [17, Chapter II, Theorems 1.8 & 1.9].
@@
Note, that while [4, Theorem 5.5.6] contains formal results, on how to derive error estimates also for $p$ in a weaker norm, e.g. $H^{-1}(\Omega)$, the argument requires $H^{2}$ regularity results for the compressible Stokes equations. As the corresponding results in [23, 9] require an additional decaying condition for the compressibility data, this makes derivation of weaker error estimates for the pressure complicated.
@@ 4.2 Temporal discretization: the discontinuous Galerkin method
$\|{\bf{u}}_{kh}\|_{L^{2}(I;H^{1}(\Omega))}+\|{\bf{u}}_{kh}\|_{L^{\infty}(I;L^{2}(\Omega))}\leq C\|{\bf{q}}\|_{L^{2}(I;H^{-1}(\Omega))+L^{1}(I;L^{2}(\Omega))}.$
@@
As (4.11) poses a square system of linear equations in finite dimensions, it suffices to show uniqueness. This is a standard argument, making use of a discrete Gronwall Lemma, see e.g. [34, Theorem 4.13] for a proof focusing especially on the $L^{1}(I;L^{2}(\Omega)^{d})$ right hand side case. ∎
@@
In our recent paper [3], we have established a best approximation type error estimate for the Stokes problem in the $L^{\infty}(I;L^{2}(\Omega)^{d})$ norm. From this more general result, we obtain in the case of homogeneous initial data the following result, see [3, Corollary 6.4].
@@
Using the error estimates for the stationary Stokes Ritz projection of Theorem 4.1, in [3, Theorem 7.4] the following estimate in terms of explicit orders of convergence was shown.
@@
The above results are valid in $L^{2}(I;L^{2}(\Omega)^{d})$ norm as well. However, using the energy and duality arguments it is possible to show the corresponding results log-free and with less regularity assumptions on the data (cf. [24, Theorem 11 & Theorem 13]).
@@
$\|{\bf{u}}-{\bf{u}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}\leq C(h^{2}+k)\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf{q}}}\|_{L^{2}(I;L^{2}(\Omega))}.$
@@ 5 Variational Discretization of the Optimal Control Problem
In this section we consider the optimal control problem subject to the fully discretized Stokes equations. We consider a variational discretization for the controls, i.e., do not fix a finite dimensional approximation of the control space yet. The problem reads
Minimize $\text{Minimize }\;J({\bf{q}}_{kh},{\bf{u}}_{kh})=\frac{1}{2}\lVert{\bf{u}}_{kh}-{\bf{u}}_{d}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\frac{\alpha}{2}\lVert{\bf{q}}_{kh}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}$ (5.1a)
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}G_{\bf{w}}(\bar{\bf{u}}_{kh})|_{I_{m}}\leq\beta,\ m=1,...,M},\quad\bar{\mu}_{kh}\geq 0\quad\text{and}\quad\langle\bar{\mu}_{kh},b-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}G_{\bf{w}}(\bar{\bf{u}}_{kh})}\rangle=0;$ (5.5b)
@@
$\|\bar{\bf{q}}_{kh}\|_{L^{\infty}(I;L^{2}(\Omega))}+\|\bar{{\bf{u}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}+\|\bar{\mu}_{kh}\|_{L^{1}(\bar{I})}\leq C.$
@@
$J(\bar{{\bf{q}}}_{kh},\bar{{\bf{u}}}_{kh})\leq J(\tilde{{\bf{q}}},{\bf{u}}_{kh}(\tilde{{\bf{q}}}))$ $=\frac{1}{2}\lVert{\bf{u}}_{kh}(\tilde{{\bf{q}}})-{\bf{u}}_{d}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\frac{\alpha}{2}\lVert\tilde{{\bf{q}}}\rVert^{2}_{L^{2}(I)}\leq C(T,\tilde{{\bf{q}}}),$
@@
$\|\bar{{\bf{q}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}+\|\bar{{\bf{u}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}\leq C.$ (5.7)
@@
$0\leq C+\langle\bar{\mu}_{kh},b-G_{\bf{w}}(\bar{{\bf{u}}}_{kh})\rangle-\frac{\delta}{4}\langle\bar{\mu}_{kh},1\rangle=C-\frac{\delta}{4}\langle\bar{\mu}_{kh},1\rangle.$
@@
$\sqrt{\alpha}\lVert\bar{{\bf{q}}}-\bar{{\bf{q}}}_{kh}\rVert_{L^{2}(I;L^{2}(\Omega))}+\lVert\bar{{\bf{u}}}-\bar{{\bf{u}}}_{kh}\rVert_{L^{2}(I;L^{2}(\Omega))}\leq C\,\ell_{k}(k^{\frac{1}{2}}+h),\quad\ell_{k}=\ln{\frac{T}{k}}.$
@@
$\alpha\|\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}}\|^{2}_{L^{2}(I;L^{2}(\Omega))}\leq(\bar{{\bf{z}}},\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}})_{I\times\Omega}+(\bar{{\bf{z}}}_{kh},\bar{{\bf{q}}}-\bar{{\bf{q}}}_{kh})_{I\times\Omega}:=I_{1}+I_{2}.$ (5.12)
@@
$\leq$ $2\|{\bf w}\|_{L^{2}(\Omega)}\|\bar{{\bf{u}}}_{kh}(\bar{{\bf{q}}}_{kh})-{\bf{u}}(\bar{{\bf{q}}}_{kh})\|_{L^{\infty}(I;L^{2}(\Omega))}\|\bar{\mu}\|_{C(\bar{I})^{*}}$
@@
$\leq$ $2\|{\bf w}\|_{L^{2}(\Omega)}\|\bar{{\bf{u}}}(\bar{{\bf{q}}})-({\bf{u}}_{kh}(\bar{{\bf{q}}})\|_{L^{\infty}(I;L^{2}(\Omega))}\|\bar{\mu}_{kh}\|_{L^{1}(\bar{I})}$
@@
$=$ $-\lVert\bar{{\bf{u}}}-\bar{{\bf{u}}}_{kh}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+(\bar{{\bf{u}}}-{\bf{u}}_{d},{\bf{u}}(\bar{{\bf{q}}}_{kh})-\bar{{\bf{u}}}_{kh})_{I\times\Omega}+(\bar{{\bf{u}}}_{kh}-{\bf{u}}_{d},{\bf{u}}_{kh}(\bar{{\bf{q}}})-\bar{{\bf{u}}})_{I\times\Omega}$
@@
$\alpha$ $\lVert\bar{{\bf{q}}}-\bar{{\bf{q}}}_{kh}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\lVert\bar{{\bf{u}}}-\bar{{\bf{u}}}_{kh}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}$
@@
$\leq C\ell^{2}_{k}(k+h^{2})+\|\bar{{\bf{u}}}-{\bf{u}}_{d}\|_{L^{2}(I;L^{2}(\Omega))}\|{\bf{u}}(\bar{{\bf{q}}}_{kh})-\bar{{\bf{u}}}_{kh}(\bar{{\bf{q}}}_{kh})\|_{L^{2}(I;L^{2}(\Omega))}$
$\quad+\|\bar{{\bf{u}}}_{kh}-{\bf{u}}_{d}\|_{L^{2}(I;L^{2}(\Omega))}\|{\bf{u}}_{kh}(\bar{{\bf{q}}})-\bar{{\bf{u}}}(\bar{{\bf{q}}})\|_{L^{2}(I;L^{2}(\Omega))}$
$\leq C\ell^{2}_{k}(k+h^{2})\left[1+\left(\|\bar{{\bf{u}}}\|_{L^{2}(I;L^{2}(\Omega))}+\|\bar{{\bf{u}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}+\|{\bf{u}}_{d}\|_{L^{2}(I;L^{2}(\Omega))}\right)\left(\|\bar{{\bf{q}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}+\|\bar{{\bf{q}}}\|_{L^{2}(I;L^{2}(\Omega))}\right)\right]$
@@
where in the last step, we used the boundedness of $\|\bar{{\bf{u}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}$, $\|\bar{{\bf{u}}}\|_{L^{2}(I;L^{2}(\Omega))}$, $\|\bar{{\bf{q}}}\|_{L^{2}(I;L^{2}(\Omega))}$, and $\|\bar{{\bf{q}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}$ from Theorem 3.8 and Lemma 5.7. ∎

### v2

@@ 4.1 Spatial discretization
This assumption is valid, for example, for Taylor-Hood and MINI finite elements on shape regular meshes, see [4, Assumption 7.2]. We define the space of discretely divergence-free vector fields $\mathbf{V}_{h}$ as
@@
Note that, if ${\bf{u}}\in\mathbf{V}$, then it holds $R_{h}^{S}({\bf{u}},p)\in\mathbf{V}_{h}$. Further, the Stokes Ritz projection satisfies the following stability, see [5, Theorem 5.2.1].
@@
Let us recall the following error estimates for the stationary discrete Stokes problem: See [5, Theorem 5.25], [25, Theorems 4.21, 4.25, 4.28], [16, Theorems 53.17 & 53.19] or [18, Chapter II, Theorems 1.8 & 1.9].
@@
Note, that while [5, Theorem 5.5.6] contains formal results, on how to derive error estimates also for $p$ in a weaker norm, e.g. $H^{-1}(\Omega)$, the argument requires $H^{2}$ regularity results for the compressible Stokes equations. As the corresponding results in [26, 10] require an additional decaying condition for the compressibility data, this makes derivation of weaker error estimates for the pressure complicated.
@@ 4.2 Temporal discretization: the discontinuous Galerkin method
$\|{\bf{u}}_{kh}\|_{L^{2}(I;H^{1}(\Omega))}+\|{\bf{u}}_{kh}\|_{L^{\infty}(I;L^{2}(\Omega))}\leq C\|{\bf{q}}\|_{L^{2}(I;H^{-1}(\Omega))+L^{1}(I;L^{2}(\Omega))},$
where the constant $C>0$ is independent of $k,h$.
@@
As (4.10) poses a square system of linear equations in finite dimensions, it suffices to show uniqueness. This is a standard argument, making use of a discrete Gronwall Lemma, see e.g. [38, Theorem 4.13] for a proof focusing especially on the $L^{1}(I;L^{2}(\Omega)^{d})$ right hand side case. ∎
@@
In our recent paper [4], we have established a best approximation type error estimate for the Stokes problem in the $L^{\infty}(I;L^{2}(\Omega)^{d})$ norm. From this more general result, we obtain in the case of homogeneous initial data the following result, see [4, Corollary 6.4].
@@
Using the error estimates for the stationary Stokes Ritz projection of Theorem 4.1, in [4, Theorem 7.4] the following estimate in terms of explicit orders of convergence was shown.
@@
The above results are valid in $L^{2}(I;L^{2}(\Omega)^{d})$ norm as well. However, using the energy and duality arguments it is possible to show the corresponding results log-free and with less regularity assumptions on the data (cf. [28, Theorem 11 & Theorem 13]).
@@
$\|{\bf{u}}-{\bf{u}}_{kh}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq C(h^{2}+k)\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf{q}}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}.$
@@ 5 Variational Discretization of the Optimal Control Problem
In this section we consider the optimal control problem subject to the fully discretized Stokes equations. We consider a variational discretization for the controls, i.e., do not fix a finite dimensional approximation of the control space yet, cf., [15, 21] The problem reads
Minimize $\text{Minimize }\;J({\bf{q}}_{kh},{\bf{u}}_{kh})=\frac{1}{2}\lVert{\bf{u}}_{kh}-{\bf{u}}_{d}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\frac{\alpha}{2}\lVert{\bf{q}}_{kh}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$ (5.1a)
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}G_{\bf{w}}(\bar{\bf{u}}_{kh})|_{I_{m}}\leq\beta,\ m=1,...,M},\quad\bar{\mu}_{kh}\geq 0\quad\text{and}\quad\langle\bar{\mu}_{kh},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}G_{\bf{w}}(\bar{\bf{u}}_{kh})}\rangle=0;$ (5.5b)
@@
$\|\bar{\bf{q}}_{kh}\|_{L^{\infty}(I;L^{2}(\Omega))}+\|\bar{{\bf{u}}}_{kh}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\|\bar{\mu}_{kh}\|_{L^{1}(\bar{I})}\leq C.$
@@
$J(\bar{{\bf{q}}}_{kh},\bar{{\bf{u}}}_{kh})\leq J(\tilde{{\bf{q}}},{\bf{u}}_{kh}(\tilde{{\bf{q}}}))$ $=\frac{1}{2}\lVert{\bf{u}}_{kh}(\tilde{{\bf{q}}})-{\bf{u}}_{d}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\frac{\alpha}{2}\lVert\tilde{{\bf{q}}}\rVert^{2}_{L^{2}(I)}\leq C(T,\tilde{{\bf{q}}}),$
@@
$\|\bar{{\bf{q}}}_{kh}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\|\bar{{\bf{u}}}_{kh}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq C.$ (5.7)
@@
$0\leq C+\langle\bar{\mu}_{kh},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}-G_{\bf{w}}(\bar{{\bf{u}}}_{kh})\rangle-\frac{\delta}{4}\langle\bar{\mu}_{kh},1\rangle=C-\frac{\delta}{4}\langle\bar{\mu}_{kh},1\rangle.$
@@
$\sqrt{\alpha}\lVert\bar{{\bf{q}}}-\bar{{\bf{q}}}_{kh}\rVert_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\lVert\bar{{\bf{u}}}-\bar{{\bf{u}}}_{kh}\rVert_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq C\,\ell_{k}(k^{\frac{1}{2}}+h),\quad\ell_{k}=\ln{\frac{T}{k}}.$
@@
$\alpha\|\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}}\|^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq(\bar{{\bf{z}}},\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}})_{I\times\Omega}+(\bar{{\bf{z}}}_{kh},\bar{{\bf{q}}}-\bar{{\bf{q}}}_{kh})_{I\times\Omega}:=I_{1}+I_{2}.$ (5.12)
@@
$\leq$ $2\|{\bf w}\|_{L^{2}(\Omega)}\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf{u}}_{kh}}(\bar{{\bf{q}}}_{kh})-{\bf{u}}(\bar{{\bf{q}}}_{kh})\|_{L^{\infty}(I;L^{2}(\Omega))}\|\bar{\mu}\|_{C(\bar{I})^{*}}$
@@
$\leq$ $2\|{\bf w}\|_{L^{2}(\Omega)}\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf{u}}}(\bar{{\bf{q}}})-({\bf{u}}_{kh}(\bar{{\bf{q}}})\|_{L^{\infty}(I;L^{2}(\Omega))}\|\bar{\mu}_{kh}\|_{L^{1}(\bar{I})}$
@@
$=$ $-\lVert\bar{{\bf{u}}}-\bar{{\bf{u}}}_{kh}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+(\bar{{\bf{u}}}-{\bf{u}}_{d},{\bf{u}}(\bar{{\bf{q}}}_{kh})-\bar{{\bf{u}}}_{kh})_{I\times\Omega}+(\bar{{\bf{u}}}_{kh}-{\bf{u}}_{d},{\bf{u}}_{kh}(\bar{{\bf{q}}})-\bar{{\bf{u}}})_{I\times\Omega}$
@@
$\alpha$ $\lVert\bar{{\bf{q}}}-\bar{{\bf{q}}}_{kh}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\lVert\bar{{\bf{u}}}-\bar{{\bf{u}}}_{kh}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$
@@
$\leq C\ell^{2}_{k}(k+h^{2})+\|\bar{{\bf{u}}}-{\bf{u}}_{d}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\|{\bf{u}}(\bar{{\bf{q}}}_{kh})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf{u}}_{kh}}(\bar{{\bf{q}}}_{kh})\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$
$\quad+\|\bar{{\bf{u}}}_{kh}-{\bf{u}}_{d}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\|{\bf{u}}_{kh}(\bar{{\bf{q}}})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf{u}}}(\bar{{\bf{q}}})\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$
$\leq C\ell^{2}_{k}(k+h^{2})\left[1+\left(\|\bar{{\bf{u}}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\|\bar{{\bf{u}}}_{kh}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\|{\bf{u}}_{d}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\right)\left(\|\bar{{\bf{q}}}_{kh}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\|\bar{{\bf{q}}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\right)\right]$
@@
where in the last step, we used the boundedness of $\|\bar{{\bf{u}}}_{kh}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$, $\|\bar{{\bf{u}}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$, $\|\bar{{\bf{q}}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$, and $\|\bar{{\bf{q}}}_{kh}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$ from Theorem 3.8 and Lemma 5.7. ∎